# Investigating the Indirect Object Identification circuit in Mamba

### v1

@@ Appendix B Normalized Logit Diff
min_diff = A_logit_corrupted - B_logit_corrupted max_diff = A_logit_unpatched - B_logit_unpatched possible_range = abs(max_diff - min_diff) # prevent divide by zero possible_range[possible_range == 0] = 1 logit_diff = A_logits_patched - B_logits_patched normalized_logit_diff = (logit_diff-min_diff)/possible_range
@@ D.1 EAP
![[Uncaptioned image]](extracted/5741814/figures/eap_attr_smaol.png)
![[Uncaptioned image]](extracted/5741814/figures/adj_mat_0.1.png)
![[Uncaptioned image]](extracted/5741814/figures/eap_attr_extra_smol.png)
![[Uncaptioned image]](extracted/5741814/figures/adj_mat_0.05.png)
![[Uncaptioned image]](extracted/5741814/figures/eap_attr_extra_extra_smol.png)
![[Uncaptioned image]](extracted/5741814/figures/adj_mat_0.01.png)
@@ D.3.1 Connections from embed
Missing n1: layers 1, 5, 8, 25, 28, 29, 30, 32, 34, 36 Missing n2: 30, 31, 36 Missing n3: layers 25, 27, 29, 30, 31, 34, 36, 38 Missing n4: layers 3, 9, 25, 32, 34, 40 Missing n5: layer 9, 25, 27, 31, 34, 37, 38, 40 Missing n1-n5: 33, 41-47 TODO: Output?
@@ D.3.3 Layers that are missing names
![[Uncaptioned image]](extracted/5741814/figures/28_layer_view.png)
28 is missing n1 28 does not receive n1 from anyone outputs n1 to 39
![[Uncaptioned image]](extracted/5741814/figures/31_layer_view.png)
31 is missing n2, n3, n5 31 does not receive n2 or n3 from anyone output n2-n3 to 39 31 receives n5 from 29 outputs n5 to 34
![[Uncaptioned image]](extracted/5741814/figures/32_layer_view.png)
@@
![[Uncaptioned image]](extracted/5741814/figures/33_layer_view.png)
@@
![[Uncaptioned image]](extracted/5741814/figures/34_layer_view.png)
34 is missing n1, n4, n5 34 does not receive n1 from anyone outputs n1 to 39 34 receives n4 from 32 does not output it 34 receives n5 from 28, 31 does not output it
![[Uncaptioned image]](extracted/5741814/figures/40_layer_view.png)
40 is missing n4, n5 40 does not receive n4-n5 does not output n4-n5

### v2

@@ Appendix B Normalized Logit Diff
min_diff =
A_logit_corrupted - B_logit_corrupted
max_diff =
A_logit_unpatched - B_logit_unpatched
possible_range = abs(max_diff - min_diff)
# prevent divide by zero
possible_range[possible_range == 0] = 1
logit_diff =
A_logits_patched - B_logits_patched
normalized_logit_diff =
(logit_diff-min_diff)/possible_range
@@ D.3.1 Connections from embed
Missing n1: layers 1, 5, 8, 25, 28, 29,
30, 32, 34, 36
Missing n2: 30, 31, 36
Missing n3: layers 25, 27, 29, 30, 31,
34, 36, 38
Missing n4: layers 3, 9, 25, 32, 34, 40
Missing n5: layer 9, 25, 27, 31, 34, 37,
38, 40
Missing n1-n5: 33, 41-47
TODO: Output?
@@ D.3.3 Layers that are missing names
28 is missing n1
28 does not receive n1 from anyone
outputs n1 to 39
31 is missing n2, n3, n5
31 does not receive n2 or n3 from anyone
output n2-n3 to 39
31 receives n5 from 29
outputs n5 to 34
@@
34 is missing n1, n4, n5
34 does not receive n1 from anyone
outputs n1 to 39
34 receives n4 from 32
does not output it
34 receives n5 from 28, 31
does not output it
40 is missing n4, n5
40 does not receive n4-n5
does not output n4-n5